OpenStudy (anonymous):

Solve the equation using the Zero-Product Property. -9n(5n-5) = 0

OpenStudy (anonymous):

Each of term = 0: -9n = 0 5n -5 = 0

OpenStudy (anonymous):

the answers I have are -1/9, 1 0, 1 -1/9, -1 0, -1

OpenStudy (mertsj):

-9n=0 or 5n-5=0 n=0 or n=1

OpenStudy (anonymous):

so your saying the answer is 0,1?

OpenStudy (mertsj):

What are you saying?

OpenStudy (anonymous):

im asking what the asnwer is to -9n(5n-5) = 0 out of the multiple choice answers i gave above please

OpenStudy (mertsj):

So you want to know what n is equal to?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

-9n=0 or 5n-5=0 n=0 or n=1

OpenStudy (anonymous):

So then the official answer simplified would be 0,1 ? anwser number 3?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

ok, thank you :)

OpenStudy (mertsj):

Answer number 2

OpenStudy (anonymous):

oh yeah, i meant anser 2 not 3 haha